### v1

@@ 3.2 Arrangements Matter
To form various collage prompt, we first uniformly sample a sub-dataset that contains 100,000 images from the training set of ImageNet-1K [26]. This subset is then divided into $L$ groups, and each group contains $K$ images. By executing $p$ random shuffles of the images within each group, we generated a collection of $L\times p$ unique collage prompts. We collect two evaluation datasets with a quadrant-grid collage prompt and a nine-grid collage prompt. For the quadrant-grid collage prompt, $L$ is set to be 25,000 and $p$ is set to be 5. For the nine-grid collage prompt, $L$ is set to be 11,111 and $p$ is set to be 10. These collage prompts are then sent into GPT-4V’s API for image recognition and the accuracy $y$ of each prompt can be obtained. The final evaluation dataset $\mathcal{D}$ thus includes pairs $\{\textbf{M}_{i},y_{i}\}$ for each of the $L\times p$ prompts, providing a comprehensive basis for analyzing the effectiveness of different collage configurations in visual recognition tasks.
@@
As stated the previous subsection, collage prompt M is represented as a graph. To analyze and predict the performance of these collage prompts, we employ a Graph Convolutional Neural Network (GCN) [44], denoted as $G_{\theta}(\textbf{A},\textbf{F})$, to process the graph data and predicts the expected accuracy of the collage prompt. For the $i$-th layer in $G_{\theta}(\textbf{A},\textbf{F})$, it takes the adjacent matrix A and hidden representation matrix $\textbf{H}_{i}$ as input, then the next layer’s output can be generated as follows,
$\textbf{H}_{i+1}=\sigma(\tilde{\textbf{D}}^{-\frac{1}{2}}\tilde{\textbf{A}}\tilde{\textbf{D}}^{-\frac{1}{2}}\textbf{H}_{i}\textbf{W}^{i}),$ (1)
where $\sigma(\cdot)$ is the non-linear activation function and $\textbf{H}_{0}=\textbf{X}$, $\tilde{\textbf{A}}=\textbf{A}+\textbf{I}$ is the adjacent matrix with self-connections. $\tilde{\textbf{D}}$ is the diagonal degree matrix of $\tilde{\textbf{A}}$, and $\textbf{W}^{i}\in\mathbb{R}^{d\times d}$ is the trainable weight matrix. As the layer gets deeper, there exists graph pooling to progressively reduce the node and the remaining one node at the final layer will be sent into a MLP head to predict the accuracy. Given the evaluation dataset $\mathcal{D}=\{\textbf{M}_{i},y_{i}\}_{i=1}^{L\cdot p}$, the update of the $G_{\theta}(\textbf{A},\textbf{F})$ at $k$-th iteration can be expressed as,
@@
Figure 3: An overview of proposed method LCP. Starting with a set of images, index sets are randomly initialized, which forms multiple collage graphs. After predicting the accuracy of each collage graphs via $G_{\theta^{\ast}}$, collage graphs that achieves top-$T$ accuracy are selected for crossover and mutation operations. This iterative process continues until reaching the maximum specified iteration, and we can obtain the optimized arrangements within the collage prompting.
3.3 Learning to Collage
@@
With the trained predictor $\mathbf{G}_{\theta^{*}}$, we can estimate the accuracy for various arrangements in the collage prompt, which enables the selection of the most effective arrangement to enhance recognition performance with GPT-4V. Due to the vast number of potential arrangements, it is impractical to evaluate each one to identify the optimal arrangement. To efficiently search for the best arrangement within a maximum number of iterations, we propose Learning to Collage Prompt (LCP) to achieve effectively searching.
Genetic algorithm has been widely used for non-differentiable optimization problems [33]. Following the framework of genetic algorithm, LCP contains several stages in each iteration: initialization, evaluation, selection, crossover, and mutation. In the initialization phase, for a given set of image features $\mathbf{F}$, we randomly generate a set of position index set =$\mathcal{I}=\{{\textbf{I}}_{1},{\textbf{I}}_{2},...,{\textbf{I}}_{P}\}$ with the related adjacency matrices $\mathcal{A}=\{\mathbf{A}_{1},\mathbf{A}_{2},\ldots,\mathbf{A}_{P}\}$. These matrices represent different possible arrangements of the collage prompt. Subsequently, for each adjacency matrix $\mathbf{A}_{i}$ in $\mathcal{A}$, we predict its expected accuracy using $\hat{y}_{i}=G_{\theta^{*}}(\mathbf{A}_{i},\mathbf{F})$, resulting in a set of predicted accuracy $\mathcal{Y}=\{\hat{y}_{0},\hat{y}_{1},\ldots,\hat{y}_{P}\}$. During the selection phase, we choose a subset $\tilde{\mathcal{I}}$ of arrangements from $\mathcal{I}$ that correspond to the top-$T$ accuracy in $\mathcal{Y}$, indicating the most promising arrangements.
@@
To generate the next generation of arrangements, crossover and mutation operations are applied to the selected subset $\tilde{\mathcal{I}}$. Specifically, in the crossover process, the new position index $\widehat{\textbf{I}}$ is divided into several segments. Each segment is determined by the corresponding part of each position index $\textbf{I}\in\tilde{\mathcal{I}}$. Therefore, the new position index will be the fusion of the optimal arrangements in the last iteration. In the following mutation process, a set of position indexes will be generated based on $\widehat{\textbf{I}}$. This process iteratively continues, refining the search for an arrangement that maximizes the accuracy of collage prompt recognition by GPT-4V.
4 Experiment
4.1 Datasets
We evaluate the performance of collage prompt on ImageNet-1K and 10 other common downstream image recognition datasets. Table 1 presents statistics regarding the number of test samples and label tokens for each dataset. Label tokens represent the number of tokens encoded by the GPT-4 tokenizer [^2] https://platform.openai.com/tokenizer, providing a measure of the textual labels’ complexity for each dataset.
Datasets Classes Samples Label Tokens
ImageNet-1K [26] 1,000 50,000 4,834
Caltech101 [11] 100 2,465 428
OxfordPets [22] 37 3,669 203
StanfordCars [15] 106 8,041 1,814
Flowers102 [20] 102 2,463 562
Food101 [2] 101 30,300 513
FGVCAircraft [19] 100 3,333 565
SUN397 [40] 397 19,850 1,859
DTD [7] 47 1,692 211
EuroSAT [13] 10 8,100 49
UCF101 [28] 101 3,783 526
4.1.1 Evaluation Dataset.
The evaluation dataset introduced in Section 3.2.1 is used for training the collage predictor and is derived from the ImageNet-1K training dataset, where we randomly and uniformly selected 100,000 images across all categories to construct $2\times 2$ and $3\times 3$ collage prompt evaluation datasets. We collected two groups of evaluation datasets with GPT-4V. The first group is the $2\times 2$ grid dataset, which contains 25,000 sets of $2\times 2$ grids. Each set of $2\times 2$ grids has approximately 5 different arrangements, resulting in 110,250 valid training data points with GPT-4V. The second group is the $3\times 3$ grid dataset, which contains 11,111 sets of $3\times 3$ grids. Each set of $3\times 3$ grids has about 10 different arrangements, yielding 102,646 valid training data points with GPT-4V. We will release the evaluation dataset after the acceptance of this paper.
4.2 Implementation Details
4.2.1 GPT-4V API.
We utilize the GPT-4V API [^3] https://platform.openai.com/docs/guides/vision, where the image input size can be selected from low, high, and auto options. We choose to use the low resolution, which defaults to 512$\times$512 pixels, costing only 85 tokens per image. All our images are resized to 512$\times$512 before being input into GPT-4V, thus avoiding any further resizing. Through experimental testing and cost analysis, we find that using high resolution was not cost-effective. It only offers minimal improvements in accuracy at a cost several times higher than low resolution. We recommend using low resolution for extensive GPT-4V applications. For the text prompt in GPT-4V, we follow the design of GPT4Vis [38]. We also observe that not providing the model with category labels, and relying solely on the model’s pre-learned knowledge for prediction, could result in the prediction of non-existent categories, with inconsistent category names. However, datasets with a large number of categories consume a significant amount of tokens. For instance, the category labels of ImageNet-1K require 4,834 tokens, leaving only 211 tokens for the remaining text, while images consume just 85 tokens.
4.3 Metrics
We use cost (US Dollars) and accuracy as our primary evaluation metrics and introduce two new metrics based on these two fundamental indicators.
  • Cost-Effective Ratio (CER) is calculated by dividing accuracy by cost, which can be formulated as,
    $\textbf{CER}=\frac{Grid\;Accuracy}{Cost},\emph{e.g}.\hbox{}(1\times 1\,Acc.)/(1\times 1\,Cost).$ (3)
    A higher score indicates achieving higher accuracy or lower cost. This metric aids in determining the value of utilizing an $n\times n$ grid and deciding which type of $n\times n$ grid to use.
  • Precision-Cost Efficiency (PCE) signifies the money saved per accuracy loss. Its formula is,
    $\textbf{PCE}=\frac{Cost\;Saved}{Accuracy\;Lost}=\frac{n\times n\,Cost-1\times 1\,Cost}{n\times n\,Acc.-1\times 1\,Acc.}$ (4)
    This metric directly reflects the amount of money saved for each point of accuracy lost, balancing accuracy and cost to help choose which type of $n\times n$ grid to use. The goal is to minimize accuracy loss while saving more on costs.
4.4 Results of Collage Prompting
Table 2: Cost and top-1 accuracy of using collage prompt in GPT-4V’s image recognition. Cost is the overall expense for evaluating 1,000 images. ‘R’ denotes using random arrangements within collage prompt.
Grid Size 1$\times$ 1 2$\times$2 3$\times$3
Cost Acc. Cost Acc.(R) Acc.(LCP) Cost Acc.(R) Acc.(LCP)
ImageNet-1K 51.30 62.0% 12.83 39.4% 45.7% 5.70 28.1% 33.7%
Caltech101 7.24 95.5% 1.81 88.4% 90.8% 0.80 79.1% 85.4%
OxfordPets 4.99 92.6% 1.25 70.1% 71.8% 0.55 53.2% 59.5%
StanfordCars 21.0 58.3% 5.28 29.8% 32.0% 2.34 11.6% 14.7%
Flowers102 8.58 70.6% 2.15 49.8% 51.1% 0.95 38.4% 43.8%
Food101 8.09 80.1% 2.02 62.2% 64.2% 0.90 39.9% 46.8%
Aircraft 8.61 36.0% 2.15 18.5% 17.7% 0.96 7.0% 10.3%
SUN397 21.55 57.7% 5.39 46.5% 48.7% 2.39 27.6% 36.3%
DTD 5.07 59.1% 1.27 48.6% 52.0% 0.56 37.5% 44.1%
EuroSAT 3.45 36.2% 0.86 42.9% 53.4% 0.38 30.4% 39.7%
UCF101 8.22 51.6% 2.06 55.2% 58.1% 0.91 37.9% 44.0%
Our analysis reveals that utilizing Collage Prompting with GPT-4V for image recognition significantly reduces inference costs without substantially compromising accuracy as shown in Table 2. By employing low-resolution settings and optimizing image arrangements within collages, we demonstrated that larger grid sizes ($2\times 2$ and $3\times 3$) considerably decrease usage costs—approximately to 1/4 and 1/9 of the cost for single images, respectively. Despite a decrease in Top-1 accuracy as grid size increases, our optimized arrangements for these collages effectively minimize accuracy loss, achieving up to 10% higher accuracy over random arrangements across various datasets. This approach highlights the balance between cost efficiency and performance preservation, making it a practical solution for leveraging large multi-modal models like GPT-4V in resource-constrained scenarios.
Table 3: Cost-Effective Ratio (CER) of collage prompt in GPT-4V’s image recognition across diverse datasets.
Grid Size 1$\times$1 2$\times$2 3$\times$3
CER CER(R) CER(LCP) CER(R) CER(LCP)
ImageNet-1K 1.21 3.07 3.56 4.93 5.91
Caltech101 13.19 48.84 50.17 98.33 106.16
OxfordPets 18.56 56.19 57.56 95.95 107.31
StanfordCars 2.76 5.65 6.07 4.95 6.27
Flowers102 8.23 23.22 23.82 40.28 45.94
Food101 9.90 30.75 31.74 44.39 52.06
Aircraft 4.18 8.59 8.22 7.32 10.77
SUN397 2.68 8.63 9.04 11.53 15.16
DTD 11.66 38.34 41.03 66.57 78.28
EuroSAT 10.49 49.74 61.91 79.30 103.57
UCF101 9.93 26.86 28.27 41.50 48.18
Table 4: Precision-Cost Efficiency (PCE) of collage prompt in GPT-4V’s image recognition across diverse datasets.
Grid Size 2$\times$ 2 3$\times$ 3
PCE(R) PCE(LCP) PCE(R) PCE(LCP)
ImageNet-1K 1.70 2.36 1.35 1.61
Caltech101 0.76 1.16 0.39 0.64
OxfordPets 0.17 0.18 0.11 0.13
StanfordCars 0.56 0.60 0.40 0.43
Flowers102 0.31 0.33 0.24 0.28
Food101 0.34 0.38 0.18 0.22
Aircraft 0.37 0.35 0.26 0.30
SUN397 1.44 1.80 0.64 0.90
DTD 0.36 0.54 0.21 0.30
EuroSAT -0.39 -0.15 0.53 -0.88
UCF101 0.23 0.26 0.17 0.19
Our analysis demonstrates the effectiveness of Collage Prompting in enhancing cost-efficiency across various datasets, with $n\times n$ grid collages significantly outperforming single $1\times 1$ images in terms of the Cost-Effective Ratio (CER) as shown in Table 3. The $3\times 3$ grids, optimized through our collage graph optimization method, show the most notable improvements in cost efficiency, especially in datasets with simpler, easily recognizable images. Additionally, our Precision-Cost Efficiency (PCE) analysis underscores the trade-off between cost savings and accuracy loss, highlighting that our optimized $2\times 2$ and $3\times 3$ grid arrangements achieve substantial cost savings while minimizing accuracy loss as shown in Table 4, thereby offering a balanced approach to cost-efficient image recognition with GPT-4V. Unexpected findings, such as the higher-than-anticipated accuracies for $2\times 2$ and $3\times 3$ grids in certain datasets, point to intriguing avenues for future research. Overall, these results underscore the practicality and efficiency of collage prompting in leveraging large multi-modal models for image recognition tasks under budget constraints.
4.4.1 Cost Analysis and Comparison
We estimated the training and inference costs on the ImageNet-1k dataset for CNN and ViT models using AWS cloud servers. We compared the performance of CNNs, ViTs, and collage prompting with GPT-4V in terms of accuracy, cost, and Cost-Effective Ratio (CER). As shown in Table 5, the total costs for training ResNet-50 and ViT-B/16 from scratch amounted to \$2,091.63 and \$3,909.46, respectively. Conversely, employing GPT-4V for single $1\times 1$ image prediction incurred only \$641.25. Remarkably, leveraging collage prompting with $2\times 2$ or $3\times 3$ grid configurations substantially reduced the costs to \$170.27 and \$87.75, respectively.
In terms of the Cost-Effective Ratio (CER), the disparity between collage prompting and traditional CNN and ViT models is stark. For instance, the CER for the $3\times 3$ grid is 17 times higher than that of ViT-B/16, underscoring the cost efficiency of collage prompting compared to conventional training approaches.
@@ 4.4.1 Cost Analysis and Comparison
Datasets Epochs Accuracy Training Cost Test Cost Total Cost CER

### v2

@@ 4.1 Dataset
To construct a collage prompting dataset with various arrangements, we generate different collages (i.e., A) for the same set of images (i.e., X). We first uniformly sample a sub-dataset that contains 100,000 images from the training set of ImageNet-1K. This subset is then divided into $L$ groups, and each group contains $K$ images. By executing $p$ random shuffles of the images within each group, we generated a collection of $L\times p$ unique collage prompts. We collect two collage prompting datasets with a quadrant-grid collage prompt and a nine-grid collage prompt. For the quadrant-grid collage prompt, $L$ is set to be 25,000 and $p$ is set to be 5. For the nine-grid collage prompt, $L$ is set to be 11,111 and $p$ is set to be 10. These collage prompts are then sent into GPT-4V’s API for image recognition and the accuracy $y$ of each prompt can be obtained. The final dataset $\mathcal{D}$ thus includes pairs $\{\textbf{M}_{i},y_{i}\}$ for each of the $L\times p$ prompts, providing a comprehensive basis for analyzing the effectiveness of different collage configurations in GPT-4V’s visual recognition. We conducted an in-depth analysis of the collected collage prompting datasets and observed the following patterns.
Observation 1 (Position Accuracy Variance)
Different positions within the collage grid have varying accuracy in GPT-4V’s visual recognition.
As shown in Figure 4, the top-left position in both $2\times 2$ and $3\times 3$ grids tends to have the highest accuracy, with accuracy decreasing towards the center and bottom-left positions, which have the lowest accuracy. Accuracy then improves slightly for the last row. This pattern suggests potential model fatigue when processing central images in the collage, leading to lower accuracy that recovers as the model approaches the final row. Based on this observation, a natural idea to optimize the arrangement is to place ‘hard’ images into positions with higher accuracy while leaving ‘easy’ images to remaining positions.
Observation 2 (Category Clustering)
Placing images of the same class together in a collage improves accuracy, while pushing images of the same class away from each other degrades the accuracy.
We observed that in both $2\times 2$ and $3\times 3$ collages, placing images of the same class together significantly improves GPT-4V’s overall recognition accuracy. Conversely, when the order is shuffled and images of the same class are not adjacent, the accuracy decreases. As illustrated in Figure 3(a) and 3(b), GPT-4V predicts one of the stingrays incorrectly when the great white shark and stingray are on separate diagonals in the collage, and correctly if the same class is adjacent. This improvement can be due to clustering images of the same class reduces the complexity of batch recognition for GPT-4V.
Observation 3 (Localization Errors)
GPT-4V often makes localization errors, predicting labels for adjacent images incorrectly.
We analyzed the prediction errors in $2\times 2$ and $3\times 3$ collages and found that the incorrectly predicted labels often correspond to images in adjacent positions. This indicates that the model correctly identifies the images but outputs the predictions to the wrong locations due to localization inaccuracies. For instance, in Figure 3(c) and 3(d), GPT-4V predicts the ‘automated teller’ machine and ‘red-breasted merganser’, but outputs to the wrong positions in the collage. When the arrangement order is changed, the model outputs the correctly predicted labels to the correct positions.
Figure 5: An overview of baseline method LCP. Starting with a set of images, index sets are randomly initialized, which forms multiple collage graphs. After predicting the accuracy of each collage graph via $G_{\theta^{\ast}}$, collage graphs that achieve top-$T$ accuracy are selected for crossover and mutation operations. This iterative process continues until reaching the maximum specified iteration and we can obtain the optimized arrangements.
Visual Bias Analysis Our analysis of collage prompting reveals that GPT-4V exhibits biases in visual recognition based on image placement. Specifically, (1) Position Accuracy Variance accuracy varies across collage positions, with top-left performing best and central positions showing lower accuracy, likely due to model fatigue. (2) Category Clustering grouping similar images enhances accuracy, while separating them reduces it. (3) Localization Errors labels are often misassigned to adjacent images, indicating spatial misalignment challenges. These insights highlight the need for optimized arrangements to mitigate biases. Our baseline algorithm leverages these biases to enhance collage layouts, improving recognition accuracy while maintaining cost efficiency. For further details and visual examples, see Appendix D.
4.2 Baseline: Learning to Collage
The above observations from the collage prompting dataset further highlights the importance of arrangement optimization within the collage prompt. To solve this task, we propose a baseline method Learning to Collage (LCP) in the benchmark. This baseline method involves two processes: training a collage predictor for accuracy prediction and refining collage’s arrangement via genetic algorithm.
@@ 4.2 Baseline: Learning to Collage
Training of Collage Predictor.  As stated in the previous subsection, collage prompt M is represented as a graph. To predict the performance of these collage prompts, we employ GCN Zhang et al. (2019), denoted as $G_{\theta}(\textbf{A},\textbf{F})$, to process the graph data and predicts the expected accuracy of the collage prompt.
Given the evaluation dataset $\mathcal{D}=\{\textbf{M}_{i},y_{i}\}_{i=1}^{L\cdot p}$, the update of the $G_{\theta}(\textbf{A},\textbf{F})$ at $k$-th iteration can be expressed as,
@@
Table 1: Benchmark results of collage prompting with different sizes.
Datasets Cost(\$/1k) Top-1 Accuracy CER PCE
Random Baseline Random Baseline Random Baseline
2$\times$2 ImageNet-1K \$12.83 39.4% 45.7% 4.99 5.38 5.49 10.60
Caltech101 \$1.81 88.4% 90.8% 13.37 13.52 2.15 3.18
OxfordPets \$1.25 70.1% 71.8% 13.20 13.34 1.18 1.20
StanfordCars \$5.28 29.8% 32.0% 5.65 5.88 1.74 1.83
Flowers102 \$2.15 49.8% 51.1% 9.75 9.88 1.36 1.39
Food101 \$2.02 62.2% 64.2% 11.05 11.21 1.40 1.46
Aircraft \$2.15 18.5% 17.7% 5.57 5.42 1.45 1.42
SUN397 \$5.39 46.5% 48.7% 7.12 7.29 4.23 6.02
DTD \$1.27 48.6% 52.0% 11.02 11.40 1.44 1.71
EuroSAT \$0.86 42.9% 53.4% 11.21 12.52 1.47 1.16
UCF101 \$2.06 55.2% 58.1% 10.39 10.65 1.26 1.30
3$\times$3 ImageNet-1K \$5.70 28.1% 33.7% 5.33 5.90 3.84 5.01
Caltech101 \$0.80 79.1% 85.4% 15.26 15.79 1.48 1.89
OxfordPets \$0.55 53.2% 59.5% 13.45 14.20 1.12 1.14
StanfordCars \$2.34 11.6% 14.7% 3.96 4.68 1.49 1.54
Flowers102 \$0.95 38.4% 43.8% 10.38 11.12 1.27 1.33
Food101 \$0.90 39.9% 46.8% 10.71 11.63 1.20 1.24
Aircraft \$0.96 7.0% 10.3% 3.32 4.52 1.30 1.35
SUN397 \$2.39 27.6% 36.3% 6.89 8.03 1.89 2.45
DTD \$0.56 37.5% 44.1% 11.22 12.21 1.23 1.35
EuroSAT \$0.38 30.4% 39.7% 10.50 12.14 1.70 2.40
UCF101 \$0.91 37.9% 44.0% 10.39 11.24 1.18 1.21
Table 2: Results of GPT-4V’s zero-shot visual recognition in 11 various datasetsWu et al. (2023c).
Dataset Cost(\$/1k) Top-1 Acc. CER
ImageNet-1K \$51.30 62.0% 4.30
Caltech101 \$7.24 95.5% 9.08
OxfordPets \$4.99 92.6% 10.09
StanfordCars \$21.10 58.3% 5.26
Flowers102 \$8.58 70.6% 7.52
Food101 \$8.09 80.1% 8.12
Aircraft \$8.61 36.0% 5.37
SUN397 \$21.55 57.7% 5.20
DTD \$5.07 59.1% 8.17
EuroSAT \$3.45 36.2% 7.19
UCF101 \$8.22 81.6% 8.14
@@
Arrangement optimization.  With the trained predictor $\mathbf{G}_{\theta^{*}}$, we can estimate the accuracy for various arrangements in the collage prompt, which enables the selection of the most effective arrangement to enhance recognition performance with GPT-4V. Due to the vast number of potential arrangements, it is impractical to evaluate each one to identify the optimal arrangement. To efficiently search for the best arrangement within a maximum number of iterations, we use genetic algorithm (GA) that has been widely used for non-differentiable optimization problems Wang et al. (2018a, 2019) to achieve effectively searching. Following the idea of GA, LCP alternately evaluates the quality of arrangements in the current population and searches for the optimal collage arrangement through operations such as selection, crossover, and mutation. As shown in Figure 5, LCP consists of several key stages:
  • Initialization. In the initialization phase, for a given set of image features $\mathbf{F}$, we randomly generate a set of position index set =$\mathcal{I}=\{{\textbf{I}}_{1},{\textbf{I}}_{2},...,{\textbf{I}}_{P}\}$ with the related adjacency matrices $\mathcal{A}=\{\mathbf{A}_{1},\mathbf{A}_{2},\ldots,\mathbf{A}_{P}\}$. These matrices represent different possible arrangements of the collage prompt.
  • Evaluation. Subsequently, for each adjacency matrix $\mathbf{A}_{i}$ in $\mathcal{A}$, we predict its expected accuracy using $\hat{y}_{i}=G_{\theta^{*}}(\mathbf{A}_{i},\mathbf{F})$, resulting in a set of predicted accuracy $\mathcal{Y}=\{\hat{y}_{0},\hat{y}_{1},\ldots,\hat{y}_{P}\}$.
  • Selection. During the selection phase, we choose a subset $\tilde{\mathcal{I}}$ of arrangements from $\mathcal{I}$ that correspond to the top-$T$ accuracy in $\mathcal{Y}$, indicating the most promising arrangements, which are then preserved in the next iteration.
@@
  • Crossover & Mutation. To generate the next generation of arrangements, crossover and mutation operations are applied to the selected subset $\tilde{\mathcal{I}}$. We randomly select two arrangements from $\tilde{\mathcal{I}}$ for crossover and mutation, generating new arrangements for the next iteration. Specifically, in the crossover process, we divide two position indexes $\textbf{I}\in\tilde{\mathcal{I}}$ into segments and cross a segment between them to generate two new position indexes. We retain the new position index with higher expected accuracy. To promote diversity, we randomly select a position index from $\tilde{\mathcal{I}}$ and mutate a randomly chosen segment of the position index. This iterative process continues, refining the search for an arrangement that maximizes the accuracy of collage prompt recognition by GPT-4V.
@@
Epochs Accuracy Training Cost Test Cost Total Cost CER